# Evaluating the Robustness of Adversarial Defenses in Malware Detection Systems

### v1

@@ IV-D RQ3: Baseline performance of defenses
This subsection evaluates the performance of defense models in the absence of adversarial attacks. The defense categorized into two groups: (1) defenses without adversary detectors (DNN, AT-rFGSMk, AT-MaxMA) and (2) those with adversary detectors (DNN+, KDE, DLA, ICNN, PAD-SMA), operating in either deferred or conservative modes.
@@
As summarized in Table II, models with adversary detectors in deferred mode consistently outperform their conservative counterparts. Deferred models exclude a small percentage of challenging samples, which simplifies classification and boosts accuracy. Notably, DNN+ (Deferred) achieves the highest accuracy (98.12%) and F1 score (98.12%), excluding 4.98% of samples. Similarly, DLA (Deferred) and KDE (Deferred) achieve competitive results with F1 scores of 97.86% and 97.30%, respectively, while excluding 5.54% and 3.71% of samples.
@@
Conservative-mode models prioritize strict classification, treating all suspicious samples flagged by the detector $g$ as malicious. This approach results in elevated false positive rates (FPRs), reducing overall performance. For instance, ICNN (Conservative) has an FPR of 10.63%, compared to 3.24% in its deferred counterpart. Similarly, PAD-SMA (Conservative) shows an accuracy of 94.43%, lower than its deferred counterpart (95.62%).
Table II: Effectiveness (%) of detectors in the absence of attacks.
Defense FNR (%) FPR (%) Acc (%) F1 (%) ExS (%)
DNN 2.13 3.13 97.38 97.42 -
AT-rFGSMk 1.13 6.45 96.24 96.37 -
AT-MaxMA 0.77 9.61 94.86 95.12 -
KDE (Deferred) 2.24 3.17 97.29 97.30 3.71
KDE (Conservative) 2.13 5.53 96.19 96.29 -
DLA (Deferred) 1.53 2.61 97.91 97.86 5.54
DLA (Conservative) 1.39 4.28 97.18 97.25 -
DNN+ (Deferred) 1.54 2.23 98.12 98.12 4.98
DNN+ (Conservative) 1.45 6.22 96.19 96.32 -
ICNN (Deferred) 2.05 3.24 97.37 97.48 5.11
ICNN (Conservative) 2.00 10.63 93.73 94.05 -
PAD-SMA (Deferred) 0.86 7.86 95.62 95.73 4.56
PAD-SMA (Conservative) 0.81 10.43 94.43 94.74 -
@@
Adversarially hardened models, such as AT-rFGSMk, AT-MaxMA, and PAD-SMA, exhibit lower false negative rates (FNRs), highlighting their ability to detect malicious samples effectively. However, this robustness against adversarial attacks comes at the cost of performance under benign conditions. For example, AT-rFGSMk achieves an FNR of 1.13% and an FPR of 6.45%, resulting in an accuracy of 96.24%. PAD-SMA (Deferred) achieves an FNR of 0.86% and an FPR of 7.86%, with an accuracy of 95.62%, lagging behind other deferred models due to its reliance on adversarial training, which biases its detector $g$ and increases FPR [23].
@@
Answer to RQ3, In the absence of attacks, deferred-mode defenses such as DNN+, DLA, and KDE achieve superior accuracy and F1 scores by excluding challenging samples, whereas conservative-mode models and adversarially hardened defenses face trade-offs between robustness and performance, with higher FPRs and lower accuracy when there are no adversarial attacks.
IV-E RQ4: Robustness against oblivious attacks
@@
While non-adaptive attack evaluations are necessary, they are insufficient to fully gauge robustness. Success against such weaker adversaries does not guarantee resilience to adaptive attacks [18]. However, these evaluations can still highlight vulnerabilities in defenses under minimal adversarial threats.
The robustness evaluation considers two configurations: deferred and conservative. In the deferred mode, samples flagged as suspicious by $g$ are excluded from the ASR calculation, isolating the vulnerabilities of the main classifier. In contrast, the conservative mode assumes all suspicious samples flagged by $g$ are malicious, incorporating them into the ASR calculation. Unless explicitly labeled as Deferred, results represent the conservative configuration.
@@
Table III: Robustness against Oblivious Attacks
Defense $\text{ASR}_{\infty}$ (%) ExS (%) ~$\tilde{\ell}_{0,median}$
KDE (Deferred) 100.00 8.25 4.00
KDE 91.75 - 4.00
DLA (Deferred) 100.00 5.83 4.00
DLA 94.17 - 4.00
DNN+ (Deferred) 100.00 38.63 4.00
DNN+ 61.37 - 4.00
ICNN (Deferred) 100.00 2.09 4.00
ICNN 97.91 - 4.00
PAD-SMA (Deferred) 100.00 96.68 2.00
PAD-SMA 3.32 - 2.00
Table III summarizes the evaluation results. In the deferred mode, all defenses exhibit $\text{ASR}_{\infty}$ of 100%, underscoring the inability of the malware detector to resist attacks independently. However, the conservative mode reveals substantial variations in robustness, highlighting the critical role of $g$ in mitigating adversarial threats.
Among the evaluated defenses, PAD-SMA exhibits exceptional robustness, achieving an ASR of just 3.32% in its conservative configuration. The flat ASR curve shown in Figure 3 highlights its consistent resilience across varying perturbation budgets. This superior performance, compared to ICNN—which shares a similar defense architecture—can be attributed to PAD-SMA’s incorporation of adversarial training, which significantly enhances the robustness of the adversary detector $g$.
ICNN exhibits a high ASR of 97.91%, with its steep ASR curve, as depicted in Figure 3, underscoring substantial vulnerability to small perturbations. Similarly, KDE and DLA perform poorly, with ASRs of 91.75% and 94.17%, respectively, in the conservative configuration. DNN+ shows moderate improvement, achieving an ASR of 61.37%, yet it remains significantly less robust than PAD-SMA in resisting oblivious attacks.
Answer to RQ4: PAD-SMA demonstrates unparalleled robustness against oblivious attacks, outperforming other defenses by a wide margin. The considerable vulnerability of ICNN, KDE, and DLA emphasizes the indispensable role of adversarial training in crafting effective defense mechanisms.
IV-F RQ5: Robustness Against Adaptive Attacks
@@
Table IV and Figure 4 present the obustness of each defense under adaptive attacks. DNN, KDE, DLA, DNN+, and ICNN exhibit high vulnerability, with median perturbation (~$\tilde{\ell}_{0,\text{median}}$) of 4.00 and $\text{ASR}_{10}$ exceeding 90%, indicating susceptibility to small perturbations.
In contrast, adversarially trained defenses such as PAD-SMA, AT-MaxMA, and AT-rFGSMk show significantly enhanced robustness. AT-MaxMA achieves the strongest resistance at low perturbation budgets, with $\text{ASR}_{10}=23.32\%$, outperforming PAD-SMA (36.34%) and AT-rFGSMk (51.35%). This highlights AT-MaxMA’s capability to mitigate low-budget attacks effectively.
At moderate perturbation budgets, as indicated by $\text{ASR}_{50}$, AT-MaxMA maintains its superiority, achieving the lowest value (65.17%), followed by PAD-SMA (89.79%) and AT-rFGSMk (92.01%). However, for high perturbation budgets, PAD-SMA demonstrates greater robustness, achieving the lowest $\text{ASR}_{\infty}=94.56\%$, slightly outperforming AT-MaxMA (96.62%) and other defenses. PAD-SMA’s performance reflects its ability to restrict the success of high-budget attacks, as evidenced by its flattened $\text{ASR}_{k}$ curve for $k\geq 50$.
These findings highlight a trade-off between robustness across different perturbation budgets and practical considerations. AT-MaxMA is well-suited for scenarios where small perturbations are prevalent, demonstrating superior robustness at low budgets while maintaining lower false positive rates (FPR) and higher benign classification accuracy. In contrast, PAD-SMA excels against high-budget attacks, showcasing enhanced robustness in such scenarios. However, this comes at the cost of increased FPR and computational overhead, as noted in [23]. Furthermore, PAD-SMA demonstrates significant resilience when attackers are unaware of the adversary detector $g$, as evidenced in Table III.
@@
Table IV: Robustness against Adaptive Attacks
Defense $\text{ASR}_{10}$ (%) $\text{ASR}_{50}$ (%) $\text{ASR}_{\infty}$ (%) ~$\tilde{\ell}_{0,median}$
DNN 94.75 100.00 100.00 4.00
AT-rFGSMk 51.35 92.01 99.45 10.00
AT-MaxMA 23.32 65.17 96.62 30.00
KDE 93.07 100.00 100.00 4.00
DLA 91.20 100.00 100.00 4.00
DNN+ 97.45 100.00 100.00 4.00
ICNN 94.75 100.00 100.00 4.00
PAD-SMA 36.34 89.79 94.56 13.00
@@
Answer to RQ5: Adaptive attacks expose varying degrees of robustness among defenses. DNN, KDE, DLA, DNN+, and ICNN exhibit high brittleness, with $\text{ASR}_{10}$ exceeding 90%. While AT-rFGSMk outperforms these simpler defenses, it falls short compared to AT-MaxMA and PAD-SMA. AT-MaxMA demonstrates exceptional robustness at low and moderate perturbation budgets, achieving $\text{ASR}_{10}=23.32\%$ and $\text{ASR}_{50}=65.17\%$. Conversely, PAD-SMA excels against high-budget attacks, achieving the lowest $\text{ASR}_{\infty}=94.56\%$. These results underscore the trade-offs between robustness across different perturbation levels, false positive rates, and computational demands, emphasizing the importance of aligning defenses with specific threat models and operational requirements.
V Related Work
V-A Evasion Attacks in Malware Detection
Problem-space attacks alter application components such as the manifest file, Dalvik bytecode, or inject benign artifacts to bypass detection. Android HIV [16] perturbs Dalvik bytecode to evade detectors while maintaining functionality. Pierazzi et al. [15] introduced the injection of benign bytecode ”gadgets” to fool detection models. Tang et al. [17] expanded this with DapAdv, leveraging a hierarchical attention mechanism to optimize code slicing for adversarial repackaging. Advanced techniques include reinforcement learning-based attacks like HRAT [38] and hybrid approaches using generative adversarial networks (GANs) [39] with co-evolution algorithms [10] have further improved the effectiveness of problem-space attacks. Black-box methods such as EvadeDroid [11] demonstrate success with minimal feature knowledge, relying on random perturbations in benign code.
Feature-space attacks focus on manipulating extracted feature vectors without altering the application itself. Gradient-based methods [9, 8] optimize perturbations to evade detection. Ensemble-based techniques such as  [20] combine diverse manipulation strategies for enhanced efficacy. GAN-based approaches, such as E-MalGAN [40] and the method proposed by Shahpasand et al. [41], craft adversarial examples that evade both malware detectors and adversarial detection systems.
V-B Defenses Against Evasion Attacks
Adversarial training is a well-established defense strategy that enhances model resilience by incorporating adversarial examples into the training process. While it proves effective against the perturbations encountered during training [8], this approach faces limitations, including vulnerability to novel attack styles, high computational overhead, and a tendency to overfit to specific attack methods [18].
Ensemble methods aim to enhance robustness by combining multiple classifiers or detectors. Smutz and Stavrou [19] proposed mutual agreement analysis, flagging uncertain predictions based on classifier disagreement. Ficco [21] introduced a dynamic ensemble system combining generic and specialized detectors. However, applying ensemble strategies to deep learning models remains challenging due to difficulties in coordinating diverse model predictions.
Defenses with auxiliary models utilize separate frameworks to identify adversarial examples or purify inputs. Li et al. [22] employed a Variational Autoencoder (VAE) to detect adversarial malware via reconstruction errors, while Li et al. [42] combined hash transformations with denoising autoencoders to mitigate perturbations. Despite initial success, these approaches often struggle against adaptive attacks.
Recently, hybrid approaches have gained prominence, combining multiple defense strategies to address the limitations of individual methods. For instance, PAD-SMA [23] integrates a malware detection system with an adversary detector, both strengthened through adversarial training with a diverse range of attacks. While this approach enhances robustness and generalization, it also incurs significant computational overhead and exhibits lower accuracy on benign samples, limiting its practicality in real-world scenarios.
VI Conclusions and Future Work

### v2

@@ V-D RQ2: Baseline performance of defenses
This subsection evaluates the performance of defense models in the absence of adversarial attacks. The defenses are categorized into two groups: (1) those without adversary detectors (DNN, AT-rFGSMk, AT-MaxMA) and (2) those with adversary detectors (DNN+, KDE, DLA, ICNN, PAD-SMA), operating in either deferred or conservative modes.
For clarity and consistency, we refer to conservative-mode models simply by their base names (e.g., DLA, KDE, ICNN), while the deferred-mode variants are explicitly indicated with the suffix “(Deferred)” throughout the paper.
@@
As summarized in Table V, models with adversary detectors in deferred mode consistently outperform their conservative counterparts. Deferred models exclude a small percentage of challenging samples, which simplifies classification and boosts accuracy. Notably, DNN+ (Deferred) achieves the highest accuracy (98.1%) and F1 score (98.1%), excluding 5.0% of samples. Similarly, DLA (Deferred) and KDE (Deferred) achieve competitive results with F1 scores of 97.9% and 97.3%, respectively, while excluding 5.5% and 3.7% of samples.
Table V: Effectiveness (%) of detectors in the absence of adversarial attacks.
Defense FNR FPR Acc F1 ExS
DNN 2.1 3.1 97.4 97.4 -
AT-rFGSMk 1.1 6.5 96.2 96.4 -
AT-MaxMA 0.8 9.6 94.9 95.1 -
KDE (Deferred) 2.2 3.2 97.3 97.3 3.7
KDE 2.1 5.5 96.2 96.3 -
DLA (Deferred) 1.5 2.6 97.9 97.9 5.5
DLA 1.4 4.3 97.2 97.3 -
DNN+ (Deferred) 1.5 2.2 98.1 98.1 5.0
DNN+ 1.5 6.2 96.2 96.3 -
ICNN (Deferred) 2.1 3.2 97.4 97.5 5.1
ICNN 2.0 10.6 93.7 94.1 -
PAD-SMA (Deferred) 0.9 7.9 95.6 95.7 4.6
PAD-SMA 0.8 10.4 94.4 94.7 -
@@
Conservative-mode models prioritize strict classification, treating all suspicious samples flagged by the detector $g$ as malicious. This approach results in elevated false positive rates (FPRs), reducing overall performance. For instance, ICNN has an FPR of 10.6%, compared to 3.2% in its deferred counterpart. Similarly, PAD-SMA shows an accuracy of 94.4%, lower than its deferred counterpart (95.6%).
@@
Adversarially hardened models, such as AT-rFGSMk, AT-MaxMA, and PAD-SMA, exhibit lower false negative rates (FNRs), highlighting their ability to detect malicious samples effectively. However, this robustness against adversarial attacks comes at the cost of performance under benign conditions. For example, AT-rFGSMk achieves an FNR of 1.1% and an FPR of 6.5%, resulting in an accuracy of 96.2%. PAD-SMA (Deferred) achieves an FNR of 0.9% and an FPR of 7.9%, with an accuracy of 95.6%, lagging behind other deferred models due to its reliance on adversarial training, which biases its detector $g$ and increases FPR [li2023pad].
@@
Answer to RQ2: In the absence of attacks, deferred-mode defenses such as DNN+ (Deferred), DLA (Deferred), and KDE (Deferred) achieve superior accuracy and F1 scores by excluding challenging samples, whereas conservative-mode models and adversarially hardened defenses face trade-offs between robustness and performance, with higher FPRs and lower accuracy when there are no adversarial attacks.
V-E RQ3: Robustness against oblivious attacks
@@
While non-adaptive attack evaluations are necessary, they are insufficient to fully gauge robustness. Success against these weaker adversaries does not guarantee resilience to adaptive attacks [carlini2019evaluating]. However, these evaluations can still highlight vulnerabilities in defenses under minimal adversarial threats.
The robustness evaluation considers two configurations: deferred and conservative. In the deferred mode, samples flagged as suspicious by $g$ are excluded from the ASR calculation, isolating the vulnerabilities of the main classifier. In contrast, the conservative mode assumes all suspicious samples flagged by $g$ are malicious, incorporating them into the ASR calculation. Unless explicitly labeled as Deferred, the results represent the conservative configuration.
@@
Table VI: Robustness against Oblivious Attacks
Defense $\text{ASR}_{\infty}$ ExS $\boldsymbol{\tilde{d}_{H,\mathrm{median}}}$ IQR
KDE (Deferred) 100.0 8.3 4.0 4.0
KDE 91.8 - 4.0 4.0
DLA (Deferred) 100.0 5.8 4.0 4.0
DLA 94.2 - 4.0 4.0
DNN+ (Deferred) 100.0 38.6 4.0 4.0
DNN+ 61.4 - 4.0 4.0
ICNN (Deferred) 100.0 2.1 4.0 4.0
ICNN 97.9 - 4.0 4.0
PAD-SMA (Deferred) 100.0 96.7 2.0 2.0
PAD-SMA 3.3 - 2.0 2.0
Table VI summarizes the evaluation results. In the deferred mode, all defenses exhibit $\text{ASR}_{\infty}$ of 100%, underscoring the inability of the base classifier to withstand attacks in isolation. However, the conservative mode reveals substantial variations in robustness, highlighting the critical role of the detector $g$ in mitigating adversarial threats.
Among the evaluated defenses, PAD-SMA exhibits exceptional robustness, achieving an ASR of just 3.3% in its conservative configuration. The flat ASR curve, shown in Figure 4 highlights its consistent resilience across varying perturbation budgets. This superior performance, compared to ICNN—which shares a similar defense architecture—can be attributed to PAD-SMA’s incorporation of adversarial training, which significantly enhances the robustness of the adversary detector $g$.
ICNN exhibits a high ASR of 97.9%, and its steep ASR curve, as depicted in Figure 4, underscoring substantial vulnerability to small perturbations. Similarly, KDE and DLA perform poorly, with ASRs of 91.8% and 94.2%, respectively, in the conservative configuration. DNN+ shows moderate improvement, achieving an ASR of 61.4%, yet it remains significantly less robust than PAD-SMA in resisting oblivious attacks.
Answer to RQ3: PAD-SMA demonstrates unparalleled robustness against oblivious attacks, outperforming other defenses by a wide margin. The considerable vulnerability of ICNN, KDE, and DLA emphasizes the indispensable role of adversarial training in crafting effective defense mechanisms.
V-F RQ4: Robustness Against Adaptive Attacks
@@
Table VII and Figure 5 present the robustness of each defense under adaptive attacks. DNN, KDE, DLA, DNN+, and ICNN exhibit high vulnerability, with a median perturbation ($\tilde{d}_{H,\mathrm{median}}$) of 4.0 and $\text{ASR}_{10}$ exceeding 90%, indicating susceptibility to small perturbations.
In contrast, adversarially trained defenses such as PAD-SMA, AT-MaxMA, and AT-rFGSMk show significantly enhanced robustness. AT-MaxMA achieves the strongest resistance at low perturbation budgets, with $\text{ASR}_{10}=23.3\%$, outperforming PAD-SMA (36.3%) and AT-rFGSMk (51.4%). This highlights AT-MaxMA’s capability to mitigate low-budget attacks effectively.
At moderate perturbation budgets, as indicated by $\text{ASR}_{50}$, AT-MaxMA maintains its superiority, achieving the lowest value (65.2%), followed by PAD-SMA (89.8%) and AT-rFGSMk (92.0%). However, for high perturbation budgets, PAD-SMA demonstrates greater robustness, achieving the lowest $\text{ASR}_{\infty}=94.6\%$, slightly outperforming AT-MaxMA (96.6%) and other defenses. PAD-SMA’s performance reflects its ability to restrict the success of high-budget attacks, as evidenced by its flattened $\text{ASR}_{k}$ curve for $k\geq 50$.
These findings highlight a clear trade-off between robustness at different perturbation budgets and practical deployment factors. AT-MaxMA is particularly effective against small perturbations, exhibiting superior robustness at low budgets while maintaining lower false-positive rates (FPR) and higher benign accuracy. Conversely, PAD-SMA demonstrates stronger resistance under high-budget attacks ($k\geq 50$) but at the cost of higher FPR. Its consistent performance in both oblivious and adaptive settings (see Table VI) underscores the benefits of adversarial training for comprehensive defense coverage.
@@
Table VII: Robustness against Adaptive Attacks
Defense $\text{ASR}_{10}$ $\text{ASR}_{50}$ $\text{ASR}_{\infty}$ $\boldsymbol{\tilde{d}_{H,\mathrm{median}}}$ IQR
DNN 94.8 100.0 100.0 4.0 4.0
AT-rFGSMk 51.4 92.0 99.5 10.0 20.0
AT-MaxMA 23.3 65.2 96.6 30.0 52.0
KDE 93.1 100.0 100.0 4.0 5.0
DLA 91.2 100.0 100.0 4.0 4.0
DNN+ 97.5 100.0 100.0 4.0 4.0
ICNN 94.8 100.0 100.0 4.0 3.0
PAD-SMA 36.3 89.8 94.6 13.0 21.0
@@
Answer to RQ4: Adaptive attacks expose varying degrees of robustness among defenses. DNN, KDE, DLA, DNN+, and ICNN exhibit substantial vulnerability, with $\text{ASR}_{10}$ exceeding 90%. While AT-rFGSMk outperforms these simpler defenses, it falls short compared to AT-MaxMA and PAD-SMA. AT-MaxMA demonstrates exceptional robustness at low and moderate perturbation budgets, achieving $\text{ASR}_{10}=23.3\%$ and $\text{ASR}_{50}=65.2\%$. Conversely, PAD-SMA excels against high-budget attacks, achieving the lowest $\text{ASR}_{\infty}$ (94.6%). These results underscore the trade-offs between robustness across different perturbation levels and false positive rates, emphasizing the importance of aligning defenses with specific threat models.
V-G RQ5: Trade-offs Between Robustness, Accuracy, and Computational Cost
To ensure a fair comparison across all defenses, we evaluate only the conservative configurations, in which every input is classified without sample deferral or exclusion. Deferred configurations are omitted from this analysis, as they can inflate performance by discarding ambiguous samples. This setup enables a consistent comparison between detector-based defenses (e.g., KDE, DLA, ICNN, PAD-SMA) and adversarially trained defenses (e.g., AT-rFGSMk, AT-MaxMA), all operating under identical conditions where each sample must be processed by the classifier. The computational cost figures reported in Table VIII are derived directly from the normalized training times presented in Appendix A-A (Table IX), ensuring the consistency across all evaluations.
Table VIII: Trade-offs among robustness, non-adversarial accuracy, and computational cost for conservative configurations.
Defense Non-adv. Acc. (%) Robustness ($\text{ASR}_{10}/\text{ASR}_{\infty}$) Training Cost (×DNN) Trade-off Summary
DNN 97.4 94.8 / 100.0 1.0 High accuracy, low robustness, minimal cost
AT-rFGSMk 96.2 51.4 / 99.5 2.8 Moderate robustness, efficient training
AT-MaxMA 94.9 23.3 / 96.6 8.9 Strong low-budget robustness, moderate cost
KDE 96.2 93.1 / 100.0 2.4 Accurate but weakly robust
DLA 97.2 91.2 / 100.0 1.1 Accurate, lightweight, poor robustness
DNN+ 96.2 97.5 / 100.0 1.1 Balanced accuracy, negligible robustness gain
ICNN 93.7 94.8 / 100.0 1.8 Low robustness, moderate cost
PAD-SMA 94.4 36.3 / 94.6 74.8 Best high-budget robustness, extreme cost
1) Accuracy Versus Robustness. Defenses optimized for adversarial robustness, such as PAD-SMA, AT-rFGSMk, and AT-MaxMA, exhibit strong resistance to adaptive attacks but show moderate degradation in their classification performance on non-adversarial samples. As summarized in Table VIII, PAD-SMA achieves the highest robustness ($\text{ASR}_{10}=36.3\%,\text{ASR}_{\infty}=94.6\%$), yet its non-adversarial accuracy drops to 94.4%, reflecting the classical robustness–accuracy trade-off [zhang2019theoretically]. A similar pattern is observed for AT-rFGSMk and AT-MaxMA, which deliver superior robustness relative to standard models but with a slight loss in clean performance. In contrast, defenses with adversary detectors such as KDE, DLA, and ICNN achieve non-adversarial accuracies comparable to adversarially trained models (typically 94–97%) but collapse under adaptive attacks ($\text{ASR}_{10}>90\%$). This indicates that while these models can correctly classify normal samples under clean conditions, they provide little effective robustness once the attacker is aware of and explicitly targets the detector $g$.
2) Robustness versus Computational Cost. Training complexity varies significantly across defense categories. Lightweight defenses (DNN, DLA, KDE, DNN+, ICNN) complete training within approximately $1\times$$2\times$ the baseline DNN cost. In contrast, adversarially trained defenses incur substantially higher computational overhead due to on-the-fly adversarial example generation. AT-rFGSMk requires roughly $2.8\times$ longer training time, while AT-MaxMA increases the cost nearly $9\times$ owing to its multi-attack generation process. The most computationally demanding approach, PAD-SMA, consumes approximately $75\times$ the baseline training time, as it jointly optimizes a classifier and an adversary detector while synthesizing adversarial samples via a Stepwise Mixture of Attacks. Despite this cost, PAD-SMA delivers unmatched robustness under high perturbation budgets, highlighting the inherent cost–robustness trade-off in adversarial training.
3) Multi-objective Perspective. No single defense simultaneously maximizes robustness, non-adversarial accuracy, and computational efficiency. AT-MaxMA achieves the most favorable compromise, combining strong robustness against low- and mid-budget attacks ($\text{ASR}_{10}=23.3\%,\text{ASR}_{50}=65.2\%$) with moderate training overhead ($8.9\times$). PAD-SMA achieves the best robustness at high perturbation levels but at prohibitive cost, making it suitable primarily for high-security environments. In contrast, lightweight defenses such as KDE or DLA offer high efficiency and strong performance on non-adversarial data but limited adversarial resistance, fitting low-threat or resource-constrained environments.
Answer to RQ5: The results reveal a clear three-way trade-off among robustness, non-adversarial accuracy, and computational efficiency. Adversarially trained defenses (AT-rFGSMk, AT-MaxMA, PAD-SMA) substantially enhance robustness but incur accuracy loss and higher computational cost. Among them, AT-MaxMA achieves the most practical balance, offering robust protection against small perturbations with moderate overhead. PAD-SMA attains the highest robustness at large perturbation budgets but at prohibitive cost. Detector-based defenses (KDE, DLA, DNN+, ICNN) maintain strong non-adversarial performance and low cost but fail under adaptive attacks. These findings underscore the importance of selecting defense strategies based on operational constraints and the anticipated intensity of adversarial threats.
VI Discussion
The comprehensive evaluation across RQ1–RQ5 provides an integrated perspective on the robustness landscape of binary-constrained malware detection systems. The proposed $\sigma$-binary attack consistently outperforms existing gradient-based, optimization-based, and ensemble attacks, confirming its capability to effectively operate under strict binary constraints. These results demonstrate that robustness in discrete feature spaces remains a fundamental challenge—one that even advanced adversarially trained defenses struggle to overcome.
Adversarially hardened defenses such as AT-MaxMA and PAD-SMA exhibit substantial resilience against low-budget attacks but reveal performance degradation under larger perturbation budgets or adaptive threat models. In particular, PAD-SMA demonstrates the highest robustness under strong attacks but incurs elevated false-positive rates and a heavy computational cost, emphasizing that achieving robustness often comes at the expense of practicality. Conversely, lighter defenses such as KDE, DLA, and ICNN retain better performance on normal samples but collapse under adaptive attacks.
The results further underscore the inherent three-way trade-off among robustness, accuracy, and efficiency. AT-MaxMA achieves the most balanced performance across these dimensions, while PAD-SMA offers superior robustness at the cost of extreme computational overhead. These findings emphasize the need for flexible robustness strategies—defenses that can adapt to different operational contexts, security requirements, and computational budgets. More broadly, the study reveals that adversarial robustness in binary domains is not solely a model-level property but a system-level characteristic influenced by architectural design, optimization dynamics, and training objectives.
Overall, the empirical insights from this study not only strengthen the understanding of adversarial vulnerabilities in malware detection systems but also provide a methodological foundation for evaluating and improving binary-constrained defenses across other security-critical domains.
VII Threats to Validity
This section outlines the principal validity considerations and methodological limitations of the proposed framework. It complements the experimental protocol described in Section V-B and the ablation analyses in Section V-C.
Internal validity. All experiments were performed under a unified setup using fixed random seeds, identical preprocessing pipelines, and consistent software and hardware environments. Gradient normalization was applied to mitigate numerical instability. While minor non-determinism from GPU operations or floating-point arithmetic may cause negligible variations, their impact on reported results is statistically insignificant.
Feature- and problem-space consistency. The proposed attack is designed and evaluated entirely within the binary feature space, which serves as the abstraction layer for analyzing adversarial robustness in static malware detectors. Within this space, perturbations are explicitly constrained to preserve functional consistency in the underlying problem space. To maintain semantic plausibility, we adopt the approximate inverse mapping strategy commonly used in prior studies [arp2014drebin, li2020adversarial, li2023pad], where an estimated inverse function $\tilde{\phi}^{-1}$ maps binary feature modifications to feasible edits at the application level. This study follows that established abstraction and does not reimplement or extend the executable-level mapping itself, as such validation lies outside the present research scope. Future work should examine direct problem-space realizations to empirically verify functional equivalence of adversarial examples.
External validity. Experiments are conducted on the MalScan dataset [wu2019malscan], selected for its broad temporal coverage (2011–2018) and sufficiently large number of malicious applications (15,430 samples), which enables statistically reliable robustness evaluation. For comparability with earlier research, the Drebin dataset was initially considered; however, it was excluded because its malware subset (only 5,560 samples collected between 2010 and 2012) is both limited in scale and outdated. Such a small and temporally narrow corpus restricts the diversity of attack surfaces and hinders meaningful assessment of adversarial robustness, making MalScan a more representative and comprehensive benchmark for this study. Using MalScan enhances the representativeness of the results for modern malware behaviors. Nevertheless, further validation using newer datasets—such as AndroZoo or enterprise-grade corpora—would improve generalizability across evolving threat landscapes.
Construct validity. The study employed standard, widely accepted metrics including accuracy, F1 score, false positive and false negative rates, attack success rate ($\text{ASR}_{k}$), median normalized Hamming distance ($\tilde{d}_{H,\mathrm{median}}$), and interquartile range (IQR). The complete codebase, pretrained models, and hyperparameter configurations are publicly released to ensure reproducibility and facilitate independent verification. In addition, the preprocessed version of the MalScan dataset, including standardized feature representations and train/validation/test partitions, is available upon request to support future comparative studies.
VIII Conclusions and Future Work